Tohban Report 2015-09-23
From RHESSI Wiki
|Start Date:||16 Sep 2015|
|End Date:||23 Sep 2015|
|Tohban:||Juan C. Martinez Oliveros|
|Next Tohban:||Lindsay Glesener|
|List all reports|
Solar activity was moderate this week. There are currently 6 ARs on the disk with the probability of some C and M class flares in the following days.
How many GOES flares occurred?
Flares above B, C, M, X class were 19 34 3 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 2 14 2 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 1 5 0 0
There were RHESSI flares/GOES flares 50 / 56 over the time range 16-Sep-15 23-Sep-15
The SSR fill level reached a maximum of ~40% and a minimum of about ~4% during the week.
The SSR fill level seems to be better at the end of this week, but it is unclear if the reason are the changes performed to the instrument or the decrease in solar activity or both.
The cold plate 1 and 2 are at about 135.0 K and ~137 K respectively.
The plan for this week was to raise detectors 3 and 4 slow front threshold to lower reduce the number of counts observed in the spectrum below 25 kev and 45 keV, on each detector respectively. The threshold were changed to 26.2 keV and 42.2 KeV. It seems like the change made on detector 3 was a mistake made by the tohban. For detector 4 the situation is unclear.
Several changes to the front and rear fast threshold were performed during the week. All the changes are reported in the detectors update page.
Detector 4 is not working well. During the week the detector livetime went down to near zero. The fast rates and the resets are extremely high, and unfortunately there are not more changes we can do, as the fast threshold is already at 0xFF. We performed a small test on this detector to see if we can make the livetime to increase by changing drastically the HV. During one BGS pass we lower the voltage 1000V, this change made the livetime to increase to ~38%, the slow counts increased and the reset rates decrease a lot. Then we changed the HV back to the starting value, with the livetime and slows decreasing rapidly and the resets rates increasing.
Detector 6 is also showing some bad symptoms. During the week the rear livetime went to zero and also showing a reverse trend. There are no more rear fast threshold changes we can make to reduce the rear fast rates.
|Night time data (fronts)||+- 4 minutes|
|Night time data (rears)||+- 4 minutes|
|Require extra passes?||No|
|Requirement for moving pointer?||No|